Reports of 32953
VIS (IX)
ryutaro.takahashi - 10:28 Wednesday 12 November 2025 (35556) Print this report
Comment to F3 GAS oscillation (35544)

The problem of F3 GAS LVDT in 2022 might remain (klog).

OBS (SDF)
takafumi.ushiba - 9:36 Wednesday 12 November 2025 (35554) Print this report
Comment to Changes of observation.snap during O4c (34169)

We accepted the following SDFs.

k1visommt1 (fig1):

All diffs are related to the initial alignment work, so they should be accepted.

k1omc (fig2):

All diffs are also relted to the initial alignment work, so they should be accepted.

k1visitmxt (fig3):

This change is related to the work reported in klog35544.
So, it should be accepted.

Images attached to this comment
VIS (IX)
shinji.miyoki - 8:05 Wednesday 12 November 2025 (35553) Print this report
Comment to F3 GAS oscillation (35544)

EY GASs show strange oscillations maybe originated from F1 at the level of 1000 counts with ~ 30 minutes period. This oscillation seemed to exist over 1 year or more.

Images attached to this comment
VIS (IX)
shinji.miyoki - 8:01 Wednesday 12 November 2025 (35552) Print this report
Comment to F3 GAS oscillation (35544)

IX F2 or F3 has also strange oscillation with very long period (~ 12 hours) as Fig.1 from 11days ago. If this is related with the tidal motion, why only this GAS is sensitive? Please check it. (Fig.1)

In addition, F3 still has an strange "excitation" at the level of 100 counts. (Fig.2/3)

Images attached to this comment
VIS (IX)
shinji.miyoki - 7:29 Wednesday 12 November 2025 (35551) Print this report
Comment to F3 GAS oscillation (35544)

In addition to this ~Hz level oscillation, it has also 0.04Hz (~25 seconds period) osccullation as Fig.1/2/3 just after the big earth recovery and yesterday. 

Images attached to this comment
MIF (General)
shinji.miyoki - 6:32 Wednesday 12 November 2025 (35550) Print this report
Comment to Initial Alignment (35546)

It is better to check the relation between tehse strange oscillation (1~2 Hz and 0.04Hz(25 seconds period) ) and OMC PD satulation.

MIF (General)
shinji.miyoki - 6:28 Wednesday 12 November 2025 (35549) Print this report
Comment to Initial Alignment (35546)

Around 0:30, IFO cpuld reach PRFPMI+RF. Around 6:00 I requested OBSERVATION, and IFO could reached it.

I also reduced  IM heaters current for 15W PRFPMI+DC lock.

CAL (General)
takahiro.yamamoto - 3:32 Wednesday 12 November 2025 (35548) Print this report
Comment to Dedicated DARM-Only Calibration (35540)
Following values are quick summary of calibration by free-swing (*_f) and pcal (*_p). (I haven't made related plots. I may be able to make them tomorrow.)
Percentages in [], (), and {} are deviation from last full calibration in Jun30 (klog#34423), last weekly calibration on Aug8 (klog#34798), and last quick calibration on Nov5 (klog#35499), respectively.

No significant changes haven't been found. But we have some minor questions for some measurements. So we plan to re-run some measurements for checking reproducibility and to re-measure some TFs with different measurement params to provide error budget for the open data.


H_mich_f = 9.742e+08 +/- 0.025e+08 [-13%]

A_BS_f = -6.072e-11 +/- 0.018e-11 [-1.7%]
A_BS_p = -6.089e-11 +/- 0.061e-11 [+1.5%]

A_IX(0)/A_BS = 0.04943 +/- 0.00041 [-1.4%]

A_IX(0)_f = -3.002e-12 +/- 0.027e-12 [-2.0%]
A_IX(0)_p = -3.010e-12 +/- 0.017e-12 [+1.2%]

A_IX(3)/A_IX(0) = 1.1303 +/- 0.0029 [-1.3%]

A_IX(3)_f = -3.393e-12 +/- 0.031e-12 [-1.9%]
A_IX(3)_p = -3.402e-12 +/- 0.017e-12 [+1.3%]

A_EXTM/A_IX(3) = 1.1393 +/- 0.0034 [-0.55%]

A_EXTM_f = -3.865e-12 +/- 0.038e-12 [-2.4%]
A_EXTM_p = -3.876e-12 +/- 0.015e-12 [+0.48%] (+0.53%) {-0.41%}

A_EXIM/A_EXTM = 38.7 +/- 1.4 [-8.1%]

A_EXIM_f = -1.496e-10 +/- 0.056e-10 [-10%]
A_EXIM_p = -1.593e-10 +/- 0.098e-10 [-5.1%] (+7.0%)
A_EXIM_fp = -1.500e-10 +/- 0.055e-10 [-7.6%]

H_DARM_f = 2.104e+12 +/- 0.025e+10 [-4.8%]
H_DARM_p = 2.098e+12 +/- 0.017e+10 [-7.9%] (+8.4%) {-1.6%}

1/H_DARM_f = 4.754e-13 +/- 0.058e-13 [+5.1%]
1/H_DARM_p = 4.767e-13 +/- 0.039e-13 [+8.6%]
MIF (General)
shinji.miyoki - 0:09 Wednesday 12 November 2025 (35546) Print this report
Initial Alignment

I performed the Initial Alignment from 23:00 to 24:00. After it, IFO could pass through ENGAGE PRMI 3F and reached IR handover. PRMI alignment seemed to be degraded before.

After the handover, the AS image showed split spots like TEM01. Although the ASC started, IFO was unlocked several times. IFO at present could not reach PRFPMI+RF lock state.

Comments to this report:
shinji.miyoki - 6:28 Wednesday 12 November 2025 (35549) Print this report

Around 0:30, IFO cpuld reach PRFPMI+RF. Around 6:00 I requested OBSERVATION, and IFO could reached it.

I also reduced  IM heaters current for 15W PRFPMI+DC lock.

shinji.miyoki - 6:32 Wednesday 12 November 2025 (35550) Print this report

It is better to check the relation between tehse strange oscillation (1~2 Hz and 0.04Hz(25 seconds period) ) and OMC PD satulation.

VIS (IX)
tomotada.akutsu - 23:58 Tuesday 11 November 2025 (35545) Print this report
Comment to F3 GAS oscillation (35544)

Do you have any particular reason to increasing the set point value but not decreasing it?

VIS (IX)
ryutaro.takahashi - 22:57 Tuesday 11 November 2025 (35544) Print this report
F3 GAS oscillation

ITMX was oscillating at the 1urad level in pitch. I found the LVDT signal of the F3 GAS is strange. The F3 GAS started oscillations around 1Hz when the LVDT signal crossed the setpoint. I changed the setpoint from 554 to 560. Though the LVDT signal was still a little strange, the fluctuation became smaller.

Images attached to this report
Comments to this report:
tomotada.akutsu - 23:58 Tuesday 11 November 2025 (35545) Print this report

Do you have any particular reason to increasing the set point value but not decreasing it?

shinji.miyoki - 7:29 Wednesday 12 November 2025 (35551) Print this report

In addition to this ~Hz level oscillation, it has also 0.04Hz (~25 seconds period) osccullation as Fig.1/2/3 just after the big earth recovery and yesterday. 

Images attached to this comment
shinji.miyoki - 8:01 Wednesday 12 November 2025 (35552) Print this report

IX F2 or F3 has also strange oscillation with very long period (~ 12 hours) as Fig.1 from 11days ago. If this is related with the tidal motion, why only this GAS is sensitive? Please check it. (Fig.1)

In addition, F3 still has an strange "excitation" at the level of 100 counts. (Fig.2/3)

Images attached to this comment
shinji.miyoki - 8:05 Wednesday 12 November 2025 (35553) Print this report

EY GASs show strange oscillations maybe originated from F1 at the level of 1000 counts with ~ 30 minutes period. This oscillation seemed to exist over 1 year or more.

Images attached to this comment
ryutaro.takahashi - 10:28 Wednesday 12 November 2025 (35556) Print this report

The problem of F3 GAS LVDT in 2022 might remain (klog).

CAL (General)
Shingo Hido - 19:31 Tuesday 11 November 2025 (35543) Print this report
Comment to Dedicated DARM-Only Calibration (35540)
CAL

After checking the noise around the calibration line, we updated the measurement template and performed the measurements.
We found a peak noise close to the 28.67 Hz calibration line. Considering both
(i) the requirement to keep the lines from the pcal and ETMX within ~1 Hz of each other, and
(ii) the need to avoid this peak noise, we decided not to change the calibration line frequency.
The measured transfer functions are listed below. The analysis is ongoing.

Measurements below were taken in the same lock
1147_XPCAL_DARM.xml
1147_XPCAL_DARM_one-third.xml
1147_XPCAL_DARM_one-tenth.xml
1147_DARM_OLG.xml
1147_ETMX_TM_DARM.xml
1147_ETMX_IM_DARM.xml
1147_YPCAL_DARM.xml
1147_ETMX_MN_DARM.xml

Measurements below were taken in the same lock
1147_ITMX_TM_DARM_RF_WFON.xml (ITMX WF 3 stages)

Measurements below were taken in the same lock
1147_ITMX_TM_DARM_RF_WFOFF.xml (ITMX WF 0 stages)

Measurements below were taken in the same lock
1147_ITMX_TM_DARM.xml
1147_ETMX_TM_DARM-1.xml
OBS (General)
ryutaro.takahashi - 17:08 Tuesday 11 November 2025 (35542) Print this report
Observation shift summary (Nov. 11, 2025)

Operator name: Miyoki, Takahashi

Shift time: 09-17 JST

Check Items:
VAC: No issues were found.
CRY cooler: No issues were found.
Compressor: No issues were found.
Temp check: No issues were found.
VIS GAS filter output check:  No issues were found.

IFO state (JST):

09:00 Start of shift: STANDBY

09:20~09:50 Initial alignment

10:24 Start of calibration: OBSERVATION WITHOUT LINES

17:00 End of shift: LOCKING

CAL (General)
Shingo Hido - 17:15 Monday 10 November 2025 (35540) Print this report
Dedicated DARM-Only Calibration
Hido, Sawada

We performed MICH-related transfer function measurements to estimate the optical gain from the FSM.
Since these were carried out before updating the MICH template (e.g. the frequency list), we may need to repeat the measurements depending on the analysis results.

Due to the earthquake, we have not yet checked the noise conditions at the cal line frequencies or measured DARM-related transfer functions.
Comments to this report:
Shingo Hido - 19:31 Tuesday 11 November 2025 (35543) Print this report
CAL

After checking the noise around the calibration line, we updated the measurement template and performed the measurements.
We found a peak noise close to the 28.67 Hz calibration line. Considering both
(i) the requirement to keep the lines from the pcal and ETMX within ~1 Hz of each other, and
(ii) the need to avoid this peak noise, we decided not to change the calibration line frequency.
The measured transfer functions are listed below. The analysis is ongoing.

Measurements below were taken in the same lock
1147_XPCAL_DARM.xml
1147_XPCAL_DARM_one-third.xml
1147_XPCAL_DARM_one-tenth.xml
1147_DARM_OLG.xml
1147_ETMX_TM_DARM.xml
1147_ETMX_IM_DARM.xml
1147_YPCAL_DARM.xml
1147_ETMX_MN_DARM.xml

Measurements below were taken in the same lock
1147_ITMX_TM_DARM_RF_WFON.xml (ITMX WF 3 stages)

Measurements below were taken in the same lock
1147_ITMX_TM_DARM_RF_WFOFF.xml (ITMX WF 0 stages)

Measurements below were taken in the same lock
1147_ITMX_TM_DARM.xml
1147_ETMX_TM_DARM-1.xml
takahiro.yamamoto - 3:32 Wednesday 12 November 2025 (35548) Print this report
Following values are quick summary of calibration by free-swing (*_f) and pcal (*_p). (I haven't made related plots. I may be able to make them tomorrow.)
Percentages in [], (), and {} are deviation from last full calibration in Jun30 (klog#34423), last weekly calibration on Aug8 (klog#34798), and last quick calibration on Nov5 (klog#35499), respectively.

No significant changes haven't been found. But we have some minor questions for some measurements. So we plan to re-run some measurements for checking reproducibility and to re-measure some TFs with different measurement params to provide error budget for the open data.


H_mich_f = 9.742e+08 +/- 0.025e+08 [-13%]

A_BS_f = -6.072e-11 +/- 0.018e-11 [-1.7%]
A_BS_p = -6.089e-11 +/- 0.061e-11 [+1.5%]

A_IX(0)/A_BS = 0.04943 +/- 0.00041 [-1.4%]

A_IX(0)_f = -3.002e-12 +/- 0.027e-12 [-2.0%]
A_IX(0)_p = -3.010e-12 +/- 0.017e-12 [+1.2%]

A_IX(3)/A_IX(0) = 1.1303 +/- 0.0029 [-1.3%]

A_IX(3)_f = -3.393e-12 +/- 0.031e-12 [-1.9%]
A_IX(3)_p = -3.402e-12 +/- 0.017e-12 [+1.3%]

A_EXTM/A_IX(3) = 1.1393 +/- 0.0034 [-0.55%]

A_EXTM_f = -3.865e-12 +/- 0.038e-12 [-2.4%]
A_EXTM_p = -3.876e-12 +/- 0.015e-12 [+0.48%] (+0.53%) {-0.41%}

A_EXIM/A_EXTM = 38.7 +/- 1.4 [-8.1%]

A_EXIM_f = -1.496e-10 +/- 0.056e-10 [-10%]
A_EXIM_p = -1.593e-10 +/- 0.098e-10 [-5.1%] (+7.0%)
A_EXIM_fp = -1.500e-10 +/- 0.055e-10 [-7.6%]

H_DARM_f = 2.104e+12 +/- 0.025e+10 [-4.8%]
H_DARM_p = 2.098e+12 +/- 0.017e+10 [-7.9%] (+8.4%) {-1.6%}

1/H_DARM_f = 4.754e-13 +/- 0.058e-13 [+5.1%]
1/H_DARM_p = 4.767e-13 +/- 0.039e-13 [+8.6%]
OBS (General)
ryutaro.takahashi - 17:04 Monday 10 November 2025 (35541) Print this report
Observation shift summary (Nov. 10, 2025)

Operator name: Miyoki, Takahashi

Shift time: 09-17 JST

Check Items:
VAC: No issues were found.
CRY cooler: No issues were found.
Compressor: No issues were found.
Temp check: No issues were found.
VIS GAS filter output check:  No issues were found.

IFO state (JST):
09:00 Start of shift: STANDBY
17:00 End of shift: LOCKING

OBS (General)
takashi.uchiyama - 9:17 Monday 10 November 2025 (35539) Print this report
KAGRA tour on two students
2025/10/31

Miyakawa, Watanabe, Sato, Uchiyama

Sorry for late post.

We took two students on a KAGRA tour.
We entered the KAGRA tunnel at about 11PM and left at 2PM. We took them the 1st floor of the center area and the Xarm tunnel.
OBS (VIS)
shinji.miyoki - 9:08 Monday 10 November 2025 (35537) Print this report
[O4c] VIS trip

Because ITMX was tripped due to large earthquakes around Sanriku-Oki in the Tohoku area around 17:00 last day, I requested

  • python3 /users/OBS/O4c/script/TRIP_reset.py.

TMSX trip seemed to be recovered. GRX trans shows ~ 0.5, while GRY shows ~ 1.1. So, PR3 or I/ETMSX alignment should be adjusted?

I requested DOWN for LSC_LOCK guardian.

I have also adjusted IM heater current for IY and EX to avoid too much cooling around 18:00 last day.

Images attached to this report
OBS (General)
takashi.uchiyama - 9:04 Monday 10 November 2025 (35538) Print this report
KAGRA tour of Professor Kato of Earthquake Institute
2025/11/07

Yokozawa, Washimi, Prof. Kato, Uchiyama

We took Prof. Kato of the Earthquake Research Institute on a KAGRA tour.
We entered the tunnel at about 1 p.m. and left the tunnel at 3 p.m. We showed the 1st floor of the center area first, then the geophysical interferometer at 500 m of the Xarm tunnel.

MIF (ASC)
shinji.miyoki - 15:25 Sunday 09 November 2025 (35536) Print this report
IFO cannot pass FIND IR Resonance after the successive earthquakes around Tohoku area.

IFO could not pass FIND IR RESONANCE after the successive earthquakes around Tohoku area.

GRX trans became 0.6~0.5, while GRY ~ 1.0. The max IR tans XY was ~0.6. So, PR3 alignment is not good?

Images attached to this report
OBS (Summary)
takayuki.tomaru, Kazuhiro Yamamoto - 9:03 Sunday 09 November 2025 (35534) Print this report
Operation Shift Nov. 8th
No special events occurred.
The sensitivity was about 3 Mpc, and the lock stabilization time was about 3 hours.
There were no earthquakes, and the sea seemed relatively calm.
LAS (General)
shinji.miyoki - 14:31 Saturday 08 November 2025 (35533) Print this report
Innolight MOPA for Emergency Case

I checked the Innolight MOPA 15NE (15W) at the Kashiwa campus. Anyway, it seemed to emit IR. I will check its power and frequency noise in the future. We can use its master laser part if it is healthy.

I also found one Mephisto 2000NE (maybe very old and Hasegawa-kun used bout 10 years ago. 2.3A for 2W output). I sent it to Kamioka and want to check it in the system that is being constructed in Toyama University.

Images attached to this report
CRY (General)
shinji.miyoki - 13:33 Saturday 08 November 2025 (35532) Print this report
Comment to IY, EX Im heater Adjustment (35502)

The temperatures of IY and EX IM could be stabilized, while EY IM temperature is increasing due to the heat absorption of ETMY.

Images attached to this comment
MIF (General)
takahiro.yamamoto - 21:05 Friday 07 November 2025 (35531) Print this report
The Return of the Combs

For checking that the frequency points at the swept sine of calibration measurements are contaminated by spectral peaks, I made a latest ASD with fine (~mHz) resolution.
Then I found that 4.17Hz and 9.77Hz combs came back which were seen also on OMC dark and CARM error signal, respectively as shown in Fig.1 and Fig.2.

These two plots were made with the data around 15:00 JST and 23:00 JST on Nov. 6th. Former and latter locks can be seen in Fig.3 as t = 6h-11h and t = 14h-20.5h, respectively. The lock duration of these two were ~5hr and ~6.5hr and there was no human activities in order to check the stability with disabling the whitening filter stage for the DCPD (probably for former lock and surely for latter lock in my memory).

These two combs were loud during O4a and were mitigated before before O4c (Comparison around 100Hz would be the easiest to understand). I couldn't remember how to mitigate the 4.17Hz combs on OMC (pin assignment of ZSW, cabling route around AS table, or else?). Though there was possibility that noise itself became louder than before, disabling the whitening filter makes SNR of these series peaks louder. If so, there may be no way to suppress them for now because stable IFO is unavailable with engaging whitening filter. In my memory about 9.77Hz combs on CARM, they were able to mitigated by adjusting the input gain of CMS as positive. Though I was not able to track recent activities about replacing IMC-CMS precisely, now the original common mode servo chassis were used for IMC and input gain were set as positive. So I have no idea why 9.77Hz combs came back. replacing chassis and/or re-cabling changed some situation around electrical connections...?

Images attached to this report
MIF (General)
takafumi.ushiba - 18:02 Friday 07 November 2025 (35530) Print this report
Comment to Frequent kick of suspensions (35505)

I checked the OMC DC PD signals in each lock from last night.

Figure 1 shows the OMC DC PD signals at the lock from 2025/11/6 06:03:48 UTC.
Since the speed of glitches is very fast, the signals would be enhanced by a factor of 10 if the whitening filter is off.
So, if the glitch reached around upper Y-cursor, the signals would be saturated if the whitening filter is on.
There are four glitches that exceeds the upper Y cursor, so the lockloss should be happened if the whitening filter was turned on.

Figure 2 shows the enlarged view of one glitch in fig1.
Clear ring down can be seen, so it seems due to the kick of suspensions.

For the lock from 2025/11/6 11:41:06 UTC, these is no large glitch that exceeds the threshold.
For the lock from 2025/11/6 12:22:33 UTC, these is also no large glitch that exceeds the threshold but the glitches close to te threshld can be seen as shown in fig3.

For the lock from 2025/11/6 13:55:08 UTC, there is 2 glitches that exceed the thresold.
One of them is smilar glitch as shown in fig2 but the other is different: very short glitch calmed down within 1 second (fig4).
The similar short glitch can be seen in the lock from 2025/11/6 21:48:58 UTC, which seems very clear sine gaussian noise (fig5).

So, some DC PD saturation seemed suspension kick but there might be another reason of sudden glitch in OMC DC PD.

Images attached to this comment
OBS (Summary)
tatsuki.washimi - 17:24 Friday 07 November 2025 (35529) Print this report
Operation Shift Summary (Nov. 7, 2025)

Operator: YokozaWashimi
Shift time: 9:00-17:00 JST

Check Items:
VAC: No issues were found.
CRY temp: No issues were found.
CRY cooling water: No issues were found.
Temp: No issues were found.
VIS GAS filter: No issues were found.

IFO state(JST):
9:00 : STANDBY
17:00 : CALIB NOT READY

Search Help
×

Warning

×